OpenStudy (anonymous):

Prove: lim x^3 = 8. x approaches 2

OpenStudy (anonymous):

simply plug 2 into x^3, and you will see that as x approaches 2 the limit is 8

OpenStudy (anonymous):

can u write a proof is basically what im asking using epsilon and delta

OpenStudy (anonymous):

ugh, i can........

OpenStudy (anonymous):

THanks...can u..tahts how my teacher wants it

OpenStudy (anonymous):

\[if: abs(x-2) < \delta\] \[then: abs(x^3 - 8) < \epsilon\] \[-\epsilon < (x^3 - 8) < \epsilon\] \[-\epsilon + 8 < (x^3) < \epsilon + 8\] \[-cuberoot(\epsilon + 8) < x < cuberoot(\epsilon + 8)\] \[(-cuberoot(\epsilon + 8) - 2) < x - 2 < (cuberoot(\epsilon + 8) - 2)\] \[abs(x - 2) < (cuberoot(\epsilon + 8) - 2) \] \[ set \delta = (cuberoot(\epsilon + 8) - 2)\]

OpenStudy (anonymous):

how did u introduce epsilon into the problem like from teh if to teh then statement? or why is taht true?

OpenStudy (anonymous):

The statement:  has the following precise definition. Given any real number\[\epsilon>0\] there exists another real number \[\delta>0\] so that: if:\[0<\left| x-a \right|<\delta, then \left| f(x)-L \right|<\epsilon\]